HUBBLE SPACE TELESCOPE DAILY REPORT #1136 PERIOD COVERED: 0700 EST Daily Status Report as of 131/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earthflats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. The first set had eight observations. Two were underexposed, and the other six were okay. Seven observations in the second set were received and displayed. They were all reported to be underexposed. 1.2 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain a 30-minute long, dark exposure to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The OSS monitored part of the observation. The data remains on Science Tape Recorder #1, and will be assessed later. 1.3 Completed Two Sets of WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies) The WF/PC-2 was used to make observations of galaxies NGC-4158 and NGC-4062. The observations were executed as scheduled. The first set of observations included two images, which were reported to show a nice image of a galaxy. The second set of observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5481 (Ages of the Outermost Globular Clusters) The WF/PC-2 was used to make observations of a globular cluster known as PAL-4-NORTH. The observations was received and displayed in the OSS. They were reported to be nominal. More data still remains on Science Tape Recorder #3, and will be analyzed later. 1.5 Completed WF/PC-2 5465 (WF/PC-2 Imaging of Two Galactic Cluster at Z = 1.2) The WF/PC-2 was used to make deep images of galaxy cluster 3C324 at Z=1.2. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 1 Successful : 1 Scheduled Reacquisitions : 14 Successful : 14 2.2 FHST Updates: Scheduled :3 Successful :3 2.3 Significant HSTARs: None 2.4 Operations Notes: During the past operational day, there were no reported problems in any HST subsystem. Due to the standard operating procedures in place and special operations requests that were executed, the solar eclipse had no adverse impact on the observatory. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of GHRS/1 5417 (A Search for the Physical Processes on Jupiter) 3.2 Complete WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies) 3.3 Complete WF/PC-2 5465 (WF/PC-2 Imaging of Two Galaxy Clusters at Z= 1.2) 3.4 Complete WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Monitor 1C: 4-CHIP STD in F170W) 3.5 Complete WF/PC-2 5652 (SINS: The Supernova Intensive Study Target of Opportunity: Cycle 4) 3.6 Complete WF/PC-2 5375 (High-Resolution Images of Virgo Galaxies with Kineamatically Distinct Cores) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-